At the end of the current year, the accounts receivable account has a debit balance of $725,000 and sales for the year total $8,220,000.

  1. The allowance account before adjustment has a debit balance of $9,800. Bad debt expense is estimated at 3/4 of 1% of sales.
  2. The allowance account before adjustment has a debit balance of $9,800. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $31,400.
  3. The allowance account before adjustment has a credit balance of $7,500. Bad debt expense is estimated at 1/2 of 1% of sales.
  4. The allowance account before adjustment has a credit balance of $7,500. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $62,300.

Determine the amount of the adjusting entry to provide for doubtful accounts under each of the assumptions (a through d) listed above.

a. $
b. $
c. $
d. $
Answer #1

Calculate adjusting entry amount

a) Amount = 8220000*1%*3/4 = $61650

b) Amount = 31400+9800 = 41200

c) Amount = 8220000*1%*1/2 = $41100

d) Amount = 62300-7500 = $54800